# Seed-Music: A Unified Framework for High Quality and Controlled Music Generation

### v1

@@ Abstract
We introduce Seed-Music, a suite of music generation systems capable of producing high-quality music with fine-grained style control. Our unified framework leverages both auto-regressive language modeling and diffusion approaches to support two key music creation workflows: controlled music generation and post-production editing. For controlled music generation, our system enables vocal music generation with performance controls from multi-modal inputs, including style descriptions, audio references, musical scores, and voice prompts. For post-production editing, it offers interactive tools for editing lyrics and vocal melodies directly in the generated audio.
We encourage readers to listen to demo audio examples at https://team.doubao.com/seed-music.
@@ 1 Introduction
  • Domain Complexity. Music signals are highly complex, exhibiting both short-term melodic coherence and long-term structural consistency. Unlike speech, vocal music features overlapping sounds across a wide frequency range. Singing, with its wide pitch range and expressive techniques, adds another layer of complexity. The model must simultaneously generate melodic vocals, harmonic tones, and rhythmic percussion.
  • Evaluation Difficulty. Evaluating music generation models often requires domain expertise to assess artistic quality. This includes judging the appeal of melodies, the coherence of chord progressions, the presence of idiomatic structure, and the expressiveness of vocals. Many of these aesthetic qualities are deeply influenced by cultural and regional differences. Quantifying these artistic elements in music poses a challenge.
  • Data Complexity. For generative models to learn how to condition outputs on lyrics, genre, instrumentation, and song structure, annotated music data is essential. Music annotation requires special domain knowledge. While many people can transcribe speech or label images, identifying musical elements such as chords, song sections, instruments, and genres requires a strong musical background.
  • Diverse User Segments and Needs. The needs of novice musicians differ greatly from those of professionals. A text-to-music system (Agostinelli et al., 2023; Copet et al., 2024) that generates a complete audio piece from a language prompt can be transformative for a beginner, but may offer limited value to professional producers, who typically seek more granular control over compositions and access to individual instrument stems. Even among professionals, needs differ: a guitarist might need vocal editing tools, while a vocalist might want to tweak guitar or other instrument tracks.
@@
With these challenges in mind, we highlight the versatility of Seed-Music. It supports vocal and instrumental music generation, singing voice synthesis, singing voice conversion, music editing and more. Our methodology, experiments, and solutions are designed to address diverse use cases. Rather than relying on a single modeling approach, such as auto-regression (AR) or diffusion, we propose a unified framework that adapts to the evolving workflows of musicians. Our key contributions are threefold:
  • We introduce an auto-regressive language model (LM) based approach for high-quality vocal music generation conditioned on diverse and multi-modal user inputs.
  • We present a diffusion based approach for fine-grained music audio editing.
  • We propose a novel method for zero-shot singing voice conversion, which requires only 10-second singing or speech recording from a user.
In Section 2, we give a brief literature review of music generation. In Section 3, we introduce our unified framework consisting of three parts: 1) a representation learning module that converts raw audio into a highly compressed representation, 2) a generation module that predicts the representation based on various user controls, and 3) a rendering module trained to produce the waveform from the representation. Our framework establishes three representations for different scenarios: audio tokens, symbolic tokens, and vocoder latents. The corresponding design choices will be detailed. In Section 4, we dive into how our unified framework can be configured and trained to support various music generation and editing tasks. In Section 5 and Section 6, we discuss potential applications and limitations of Seed-Music, including those related to building safe and ethical generative AI systems.
@@ 2 Literature Review
Vocal music can be defined as the combination of a vocal track and an instrumental track. Historically, full mixes of vocal music were created by generating these two tracks separately and mixing them together. With advances in ML-based generative modeling, Jukebox (Dhariwal et al., 2020) was one of the first published works to output full mixes of vocal music in an end-to-end fashion. Although prohibitively slow, Jukebox could generate vocal music that closely reflected input lyrics and specified artist and genre. Fast forward to today, many AI music creation apps enable creators to instantly generate vocal music on-demand by providing natural language prompts. The field of music generation has a long history. Music generation systems can largely be divided into two categories: symbolic domain generation and audio domain generation. A few representative works are covered in this section.
Symbolic music based systems
@@
The earliest instrumental music systems were rule-based and generated notes in limited styles (Cope, 1989). These systems generate symbolic representation, which is then rendered into audio using digital signal processing (DSP) methods. Later, rule-based approaches were replaced with data-driven approaches for music generation. Examples include FolkRNN (Sturm et al., 2016), PerformanceRNN (Simon and Oore, 2017), MusicTransformer (Huang et al., 2018), MuseNet (OpenAI, ), CoCoNet (Huang et al., 2019) and many others. Although many systems relied mainly on MIDI (Music Instrument Digital Inferfaces) data, other symbolic notation such as ABC notation, tablature and music XML (Dong et al., 2020; Sarmento et al., 2021) were also explored. The defining limitation for all these systems was the limited availability of labelled training data. MIDI transcriptions are notoriously difficult, time-consuming and expensive to procure.
Audio renderers for symbolic music based systems
@@
A symbolic system, whether rule-based or data-driven, requires an audio renderer to produce sound. A traditional audio renderer employs DSP-based synthesizers and samplers to render instrument sounds. A parallel line of research applied a data-driven approach to audio synthesis, which is often referred as “neural audio synthesis”. Examples include NSynth (Engel et al., 2017), GANSynth (Engel et al., 2019), DDSP (Engel et al., 2020), MIDI-DDSP (Wu et al., 2022), WAVERNN-based approaches (Kalchbrenner et al., 2018; Hantrakul et al., 2019). These techniques opened up new timbral and tonal possibilities compared to traditional DSP-based audio synthesis (Serra and Smith, 1990). For example, a WaveNet-based audio render can generate expressive and nuanced piano recordings from generated piano scores (Hawthorne et al., 2019). In the vocal generation domain, a Singing Voice Synthesis (SVS) can render vocal performances based on notes and phonemes generated by a symbolic system (Cook, 1996; Nishimura et al., 2016; Yi et al., 2019; Lu et al., 2020; Zhuang et al., 2021).
Language model generative models
@@
Breakthroughs in language modeling gave birth to a new way of generating musical audio. AudioLM (Borsos et al., 2023) demonstrated waveform generation could be formulated as a next-token prediction task where tokens are extracted from a neural codec such as SoundStream (Zeghidour et al., 2021). In this light, we are inspired by LM based systems in speech synthesis, including a related family of Seed Models (Anastassiou et al., 2024; Bai et al., 2024), the family of VALLE-style models (Wang et al., 2023a; Zhang et al., 2023; Chen et al., 2024a) and many others (Betker, 2023; Łajszczak et al., 2024).
In music generation, LM based modeling marks a paradigm shift. MusicLM (Agostinelli et al., 2023) could train directly on full mixes of music at unprecedented scales. These new end-to-end systems are composed of a hierarchy of modules that blur the historical separation of symbolic generation and audio synthesis. From a high level, the singing content and the music style are captured by “semantic tokens”, which are manipulated by LMs at earlier stages of the hierarchy, while instrumental timbre and high frequency details are captured by “acoustic tokens” and manipulated in later stages of the hierarchy.
@@
Additionally, an LM-based approach enables other modalities like text to be effectively transformed into music. Instead of procuring an impossible amount of paired text-music data, MusicLM circumvented this need by extracting semantic tokens from pre-trained models like MuLan (Huang et al., 2022) and Wav2Vec-BERT (Chung et al., 2021). An LM-based model then transforms these semantic tokens into target audio tokens representing music. Follow-up works usually have a similar high level recipe: a neural audio codec tokenizes the audio into discrete codes, a GPT2-like system (Vaswani et al., 2017; Radford et al., 2019; Touvron et al., 2023) performs next-token prediction of audio tokens based on multi-modal inputs that are also casted into token format, and the neural audio codec futher renders the waveform based on the predicted token sequence. These methods can also be applied to singing vocals (Li et al., 2024a, b; Huang et al., 2024).
Diffusion-based generative models
@@
An explosion of diffusion-based systems are revolutionizing the way images, video and audio are generated. We take inspiration from the vision field ((Ho et al., 2020; Song et al., 2022, 2021; Rombach et al., 2022)) and adapt these for music generation.
@@
For instrumental music generation, a variety of works successfully implemented text-to-music based on a diffusion backbone: Noise2Music (Huang et al., 2023a), Stable Audio (Evans et al., 2024a), Stable Audio Open (Evans et al., 2024b), MUSTANGO (Melechovsky et al., 2024) and MusicLDM (Chen et al., 2023; Liu et al., 2024). From a high level, instead of modeling high-dimensional waveforms directly in the diffusion process, these recipes often employ latent diffusion (Rombach et al., 2022) on normalized latents extracted from a vocoder that was trained separately. The process of denoising the intermediate representation can then be conditioned on various music annotations to support fine-grain and controlled generation. These include contour-based and downbeat conditioning in Music-ControlNet (Wu et al., 2023), chroma-based melody in Music-Gen (Copet et al., 2024) or mixtures of the above in JASCO (Tal et al., 2024).
@@
In the vision domain, diffusion-based systems have generalized to tasks beyond generation from scratch. These include in-painting, style transfer and contour-based editing (Meng et al., 2021; Couairon et al., 2022; Su et al., 2023; Kawar et al., 2023; Esser et al., 2024a). We note many exciting analogues in the vocal music domain. For example, recent singing voice synthesis (SVS) (Liu et al., 2022) and singing voice conversion (SVC) (Li et al., 2024c; Chen et al., 2024b; Yamamoto et al., 2023) approaches can be thought of as a type of controllable style transfer.
@@
Representation Learning In both LM-based and diffusion-based generative modeling, a critical component is the choice of representation for the audio signal. Methods for encoding speech, environmental sound and music exist on a spectrum ranging from semantic representations to acoustic representations. On one end, CLAP (Elizalde et al., 2022) and MuLan (Huang et al., 2022) are joint music-text representations that largely capture the semantic content of the underlying music and language description. On the other end, auto-encoder style models like WaveVAE (Peng et al., 2020) or Music2Latent (Pasini et al., 2024) compress waveform into an n-dimensional continuous latent space that largely captures acoustic content. Discrete neural audio codecs such as SoundStream (Zeghidour et al., 2021), Encodec (Défossez et al., 2022), Descript Audio Codec (Kumar et al., 2023), HiFi-Codec (Yang et al., 2023) and WavTokenizer (Ji et al., 2024) capture acoustic content as a series of quantized codebooks of increasing frequency detail. Spectrogram-derived approaches such as AudioMAE (Huang et al., 2023b) and Wav2Vec-BERT (Chung et al., 2021) are pretraining stages capturing varying mixtures of acoustic and semantic information suited to the downstream task. Some approaches operate on spectrograms directly as part of a text-based in-painting framework like VoiceBox (Le et al., 2024) or AudioBox (Vyas et al., 2023) whilst others may even use aggressively down-sampled waveform in a diffusion-based framework (Huang et al., 2023a) or VQ-VAE framework (Dhariwal et al., 2020). Depending on the task at hand, these approaches present varying levels of advantages and trade-offs. What type of representation is most suitable for music, methods to compress short term and long term features, mechanisms for encoding melodic and rhythmic domain knowledge and the trade-off between semantic and acoustic information continues to be active areas of work.
@@ 3 Method
Table 1: Comparison of intermediate representations.
@@
  • Symbolic representations, such as MIDI, ABC notation and MusicXML, are discrete and can be easily tokenized into a format compatible with LMs. Unlike audio tokens, symbolic representations are interpretable, allowing users to read and modify them directly. However, their lack of acoustic details means the system has to rely heavily on the Renderer’s ability to generate nuanced acoustic characteristics for musical performance. Training such a Renderer requires large-scale datasets of paired audio and symbolic transcriptions, which are especially scarce for vocal music.
@@ 3.2 Symbolic Token-based Pipeline
Figure 3: Overview of the pipeline using symbolic tokens as the intermediate representation. (1) Conditioned on the user prompt, the auto-regressive LM generates the symbolic tokens corresponding to a lead sheet. (2) The diffusion transformer model generates continuous vocoder latents given the symbolic tokens. (3) The vocoder then generates the high-quality 44KHz stereo audio waveform.
@@
  • To extract the symbolic features from audio for training the above system, we utilize in-house MIR models, including beat tracking (Hung et al., 2022), key and chord detection (Lu et al., 2021), structural section segmentation (Wang et al., 2022), five-instrument MIDI transcription (i.e., vocals, piano, guitar, bass, and drums) (Lu et al., 2023; Wang et al., 2024a), and singing lyrics transcription. The lead sheet tokens represent note-level details such as pitch, duration, position within a bar, vocal phonemes aligned to notes, and track-level attributes like section, instrument, key, and tempo.
  • The one-to-one mapping between lead sheet tokens and human-readable lead sheets allows users to understand, edit, and interact with the musical scores directly. We experimented with different methods to generate lead sheet token sequences: REMI-style (Huang and Yang, 2020) and xVal (Golkar et al., 2023). The REMI-style method interleaves instrument tracks into a quantized beat-based format, while xVal encodes onset and duration as continuous values. Although xVal-style encoding better follows our generative model’s end product, the music performance, more closely, we found that the REMI-style one better suited for user interaction with musicians.

### v2

@@ Abstract
We introduce Seed-Music, a suite of music generation and editing systems designed to produce high-quality music with fine-grained style control. Our unified framework leverages both auto-regressive language modeling and diffusion approaches to support two key music creation workflows: controlled music generation and post-production editing. For controlled music generation, our system enables vocal music generation with performance controls from multi-modal inputs, including lyrics, style descriptions, audio references, musical scores, and voice prompts. For post-production editing, it offers interactive tools for editing vocal lyrics, melodies, and timbres directly within an existing music audio track. We encourage readers to explore the demo audio examples at https://team.doubao.com/seed-music.
@@ 1 Introduction
  • Domain complexity: Music signals are highly complex, exhibiting both short-term melodic coherence and long-term structural consistency. Unlike speech, vocal music features overlapping sounds across a wide frequency range. Singing, with its wide pitch range and expressive techniques, adds another layer of complexity. The model must simultaneously generate melodic vocals, harmonic tones, and rhythmic percussion.
  • Evaluation difficulty: Evaluating music generation models often requires domain expertise to assess artistic quality. This includes judging the appeal of melodies, the coherence of chord progressions, the presence of idiomatic structure, and the expressiveness of vocals. Many of these aesthetic qualities are deeply influenced by cultural and regional differences. Quantifying these artistic elements in music poses a challenge.
  • Data complexity: Generative models require annotated music data to learn how to condition outputs based on lyrics, genre, instrumentation, and song structure. However, music annotations require specialized domain knowledge. While tasks like speech transcription or image labeling are accessible to many, identifying musical elements such as chords, song sections, instruments, and genres requires a strong musical background.
  • Diverse user segments and needs: The needs of novice creators differ greatly from those of professionals. A text-to-music system (Agostinelli et al., 2023; Copet et al., 2024) that generates a full audio piece from a text prompt can be transformative for a beginner, but may offer limited values to professional producers, who often seek more granular control over compositions and access to individual instrument stems. Even among professionals, needs differ: a guitarist might need vocal editing tools, while a vocalist might want to tweak guitar or other instrument tracks.
@@
With these challenges in mind, we highlight the versatility of Seed-Music. It supports vocal and instrumental music generation, singing voice synthesis, singing voice conversion, music editing and more. Our methodology, experiments, and solutions are designed to address diverse use cases. Rather than relying on a single modeling approach, such as auto-regression (AR) or diffusion, we propose a unified framework that adapts to the evolving workflows of musicians.
Our key contributions are threefold:
  • We introduce a unified framework that leverages both auto-regressive language modeling and diffusion approaches for high-quality vocal music generation conditioned on diverse and multi-modal inputs.
  • We present a diffusion-based approach that enables fine-grained editing of music audio.
  • We propose a novel zero-shot singing voice conversion method, which requires only a 10-second recording of either singing or speech from the user.
In Section 2, we give a brief literature review of music generation. In Section 3, we introduce our unified framework, which is built upon three fundamental representations: audio tokens, symbolic tokens, and vocoder latents. The corresponding pipelines and design choices will be detailed. In Section 4, we dive into how our unified framework can be configured and trained to support various music generation and editing tasks. In Section 5 and Section 6, we discuss potential applications and limitations of Seed-Music, including those related to building safe and ethical generative AI systems.
@@ 2 Literature Review
Vocal music is typically defined as the combination of a vocal track and an instrumental track. Historically, these tracks were generated separately and then mixed together to create the full mixture. With advances in machine learning based generative modeling, Jukebox (Dhariwal et al., 2020) was one of the first systems capable of generating full vocal music mixes in an end-to-end fashion. While its generation was prohibitively slow, Jukebox has demonstrated the ability to produce vocal music that aligned closely with input lyrics, as well as specified artists and genres. Today, many AI music creation platforms enable creators to instantly generate vocal music on demand using natural language prompts. The field of music generation has a long history and can be broadly categorized into two areas: symbolic domain and audio domain generations. This section covers a number of important works representing both approaches.
Symbolic music based systems.
@@
Early instrumental music generation systems were rule-based, generating notes in limited styles (Cope, 1989). These systems generated symbolic forms, which were then rendered into audio using digital signal processing (DSP) methods. Over time, rule-based approaches were replaced with data-driven ones for music generation. Examples include FolkRNN (Sturm et al., 2016), PerformanceRNN (Simon and Oore, 2017), MusicTransformer (Huang et al., 2018), MuseNet (OpenAI, ), CoCoNet (Huang et al., 2019), and others. While many of these systems primarily relied on MIDI (Music Instrument Digital Interface) data, other symbolic notations such as ABC notation, tablature, and MusicXML were also explored (Dong et al., 2020; Sarmento et al., 2021). A major limitation of these systems was the scarcity of training data, as MIDI transcriptions are notoriously difficult, time-consuming, and expensive to produce.
Audio rendering for symbolic music based systems.
@@
Whether rule-based or data-driven, symbolic music generation systems require an audio renderer to produce sounds. Traditional audio renderers use DSP-based or sample-based synthesizers to render instrumental sounds. A parallel line of research has applied data-driven approaches to audio synthesis, often referred to as “neural audio synthesis”. Notable examples include NSynth (Engel et al., 2017), GANSynth (Engel et al., 2019), DDSP (Engel et al., 2020), MIDI-DDSP (Wu et al., 2022), and WaveRNN-based approaches (Kalchbrenner et al., 2018; Hantrakul et al., 2019). These techniques have expanded the range of timbral and tonal possibilities beyond traditional approaches (Serra and Smith, 1990). For instance, WaveNet-based models can render expressive and nuanced piano performances from generated scores (Hawthorne et al., 2019). In the vocal generation domain, Singing Voice Synthesis (SVS) systems render vocal performances based on notes and lyric phonemes generated by symbolic music systems (Cook, 1996; Nishimura et al., 2016; Yi et al., 2019; Lu et al., 2020; Zhuang et al., 2021).
Language model based generative approaches.
@@
Recent advancements in language modeling have introduced a new paradigm for generating musical audio. AudioLM (Borsos et al., 2023) demonstrated that waveform generation can be framed as a next-token prediction task, where tokens are extracted from a neural codec such as SoundStream (Zeghidour et al., 2021). In this work, our approach is inspired by LM-based systems in speech synthesis, including Seed-family models (Anastassiou et al., 2024; Bai et al., 2024), VALLE-style models (Wang et al., 2023a; Zhang et al., 2023; Chen et al., 2024a), and others (Betker, 2023; Łajszczak et al., 2024).
In music generation, LM-based approaches have shown great promise. MusicLM (Agostinelli et al., 2023) introduced a framework for training directly on full music mixes at an unprecedented scale. This new end-to-end system integrates modules in a hierarchical structure that blurs the traditional distinction between symbolic music generation and audio synthesis. At a higher level, “semantic tokens” capture elements such as vocal content and musical styles, which are modeled by LMs in the earlier stages of the hierarchy. In the later stages, “acoustic tokens” are handled to characterize timbral attributes and high-frequency nuances of audio.
@@
LM-based approaches also enable the incorporation of different modalities, such as text, into music generation. MusicLM utilizes semantic tokens extracted by pre-trained models like MuLan (Huang et al., 2022) and Wav2Vec-BERT (Chung et al., 2021), reducing the need for vast amounts of paired text-music data. Based on these semantic tokens, an auto-regressive LM generates audio tokens that represent music audio. Follow-up approaches typically adhere to this pipeline: a neural audio codec tokenizes audio signals into discrete codes, a GPT-2-style LM (Vaswani et al., 2017; Radford et al., 2019; Touvron et al., 2023) predicts a sequence of audio tokens from multi-modal inputs, and a neural audio codec renders the final waveform from the predicted tokens. These methods have also been successfully applied to singing vocals (Li et al., 2024a, b; Huang et al., 2024).
Diffusion-based generative models.
@@
The emergence of diffusion-based models has revolutionized the generation of images, video, and audio. Inspired by recent advances in the vision field (Ho et al., 2020; Song et al., 2022, 2021; Rombach et al., 2022), we adapt these techniques for music generation.
@@
For instrumental music generation, several works have successfully implemented text-to-music generation based on diffusion backbones. Examples include Noise2Music (Huang et al., 2023a), Stable Audio (Evans et al., 2024a), Stable Audio Open (Evans et al., 2024b), MUSTANGO (Melechovsky et al., 2024), and MusicLDM (Chen et al., 2023; Liu et al., 2024). Rather than directly modeling raw waveforms in the diffusion process, these methods often use latent diffusion (Rombach et al., 2022) on normalized latents extracted from a separately trained vocoder. The denoising process on these intermediate representations (or latents) can be conditioned on various music attributes to enable fine-grained and controlled generation. Examples include the time-varying controls conditioning (e.g. dynamic, melody, rhythm) in Music-ControlNet (Wu et al., 2023), chroma-based melody conditioning in Music-Gen (Copet et al., 2024), and combinations of these in JASCO (Tal et al., 2024).
@@
In the vision domain, diffusion-based applications have expanded beyond pure generation to tasks like in-painting, style transfer, and contour-based editing (Meng et al., 2021; Couairon et al., 2022; Su et al., 2023; Kawar et al., 2023; Esser et al., 2024a). We see exciting similarities in the vocal music domain. Recent approaches for SVS (Liu et al., 2022) and singing voice conversion (Li et al., 2024c; Chen et al., 2024b; Yamamoto et al., 2023) can be considered as a form of controllable style transfer, much like in visual applications.
@@
Representation learning. In both LM-based and diffusion-based approaches, the choice of representation for audio signals is crucial. Methods for encoding speech, environmental sounds, and music exist on a spectrum from high-level semantics to low-level acoustic representations. On the semantic end, models like CLAP (Elizalde et al., 2022) and MuLan (Huang et al., 2022) use joint audio-text representations to capture the audio semantics of music and corresponding text descriptions. On the acoustic end, autoencoder-style models such as WaveVAE (Peng et al., 2020) and Music2Latent (Pasini et al., 2024) compress waveforms into continuous latent spaces that capture the acoustic details. Discrete neural audio codecs like SoundStream (Zeghidour et al., 2021), Encodec (Défossez et al., 2022), Descript Audio Codec (Kumar et al., 2023), HiFi-Codec (Yang et al., 2023), and WavTokenizer (Ji et al., 2024) capture acoustic information using a series of quantized codebooks with increasing levels of frequency details. Spectrogram-derived approaches such as AudioMAE (Huang et al., 2023b) and Wav2Vec-BERT (Chung et al., 2021) can be used as pre-training stages to capture a varying mixture of acoustic and semantic information suited to downstream tasks. Some methods, such as VoiceBox (Le et al., 2024) and AudioBox (Vyas et al., 2023), directly manipulate spectrograms within a text-based inpainting framework, while others use auto-regressively down-sampled waveforms in diffusion-based (Huang et al., 2023a) or VQ-VAE-based frameworks (Dhariwal et al., 2020).
Each approach presents unique advantages and limitations depending on the task. Active research continues to focus on identifying optimal representations for music audio, including compression methods that balance short-term and long-term sequences, encoding techniques that can jointly capture melodic, harmonic, and rhythmic information, and strategies for navigating the trade-offs between high-level semantic and low-level acoustic features.
@@ 3 Method
Table 1: Comparison of different intermediate representations.
@@
  • Symbolic representations, such as MIDI, ABC notation and MusicXML, are discrete and can be easily tokenized into a format compatible with LMs. Unlike audio tokens, symbolic representations are interpretable, allowing creators to read and modify them directly. However, their lack of acoustic details means the system has to rely heavily on the Renderer’s ability to generate nuanced acoustic characteristics for musical performance. Training such a Renderer requires large-scale datasets of paired audio and symbolic transcriptions, which are especially scarce for vocal music.
@@ 3.2 Symbolic Token-based Pipeline
Figure 3: Overview of the pipeline using symbolic tokens as the intermediate representation. (1) Conditioned on the user prompt, the auto-regressive LM generates the symbolic tokens corresponding to a lead sheet. (2) The diffusion transformer model generates continuous vocoder latents given the symbolic tokens. (3) The vocoder then generates the high-quality 44.1KHz stereo audio waveform.
@@
  • To extract the symbolic features from audio for training the above system, we utilize in-house Music Information Retrieval (MIR) models, including beat tracking (Hung et al., 2022), key and chord detection (Lu et al., 2021), structural section segmentation (Wang et al., 2022), five-instrument MIDI transcription (i.e., vocals, piano, guitar, bass, and drums) (Lu et al., 2023; Wang et al., 2024a), and singing lyrics transcription. The lead sheet tokens represent note-level details such as pitch, duration, position within a bar, vocal phonemes aligned to notes, and track-level attributes like section, instrument, and tempo.
  • The one-to-one mapping between lead sheet tokens and human-readable lead sheets allows creators to understand, edit, and interact with the musical scores directly. We experimented with different methods to generate lead sheet token sequences: REMI-style (Huang and Yang, 2020) and xVal (Golkar et al., 2023). The REMI-style method interleaves instrument tracks into a quantized beat-based format, while xVal encodes onset and duration as continuous values. Although xVal-style encoding better follows our generative model’s end product, the music performance, more closely, we found that the REMI-style one better suited for user interaction with musicians.